OpenStudy (anonymous):

solve the system any algebraic method. 3x-5/3y=5 -3/5x1/3y=-5

OpenStudy (anonymous):

solution or method?

OpenStudy (anonymous):

method

hero (hero):

Can you use parentheses please? I cant figure out what the fraction looks like

OpenStudy (anonymous):

it is \[3x-\frac{5}{3}y=5\] \[-\frac{3}{5}x+\frac{1}{3}y=-5\] is my guess

OpenStudy (anonymous):

so y=-5 whats x ?

OpenStudy (anonymous):

if so lines are parallel and there is no solution because they do not intersect

OpenStudy (anonymous):

okay thanks :)